# Partition Functions of Determinantal Point Processes on Polarized Kähler Manifolds

### v2

@@ 1.2. Relation to physics literature
In the physics literature, the determinantal point process on Riemann surfaces corresponds to the plasma analogy in the quantum Hall effect (QHE). For a physics background on QHE, see [Ton16]. The form of asymptotic expansion of the corresponding partition function was conjectured in [ZW06] and later proved in [Kle14, KMMW17, SY25] for a pure bulk situation. For mathematical results in the presence of an edge, see [BKS23, BKSY25] and references therein.
@@ 1.3. Statement of the main results
The following Theorem generalizes formula [Lu04, (4.4)] for all $j\geq 0$. Let $X$ be a holomorphic vector field on $M$ and $\theta_{X}$ satisfy $\iota_{X}\omega=-\bar{\partial}\theta_{X}$.
@@
Let $(M,\omega)$ be a polarized Kähler manifold. Let $X\in\mathfrak{h}(M)$ and $\theta_{X}$ be given as above. Then for all $j\geq 0$, the following identity holds :
@@ 3. Asymptotic expansion of the DPP partition function
$\gamma^{j}_{\varphi}(\psi):=\int_{M}\psi(\Delta_{\varphi}a_{j-1}(\omega_{\varphi})-a_{j}(\omega_{\varphi}))\frac{\omega_{\varphi}^{n}}{n!},$ (3.39)
@@ 4. Holomorphic invariants from Bergman kernel asymptotics and critical points
Let $(M,\omega)$ be a polarized Kähler manifold. Let $X\in\mathfrak{h}(M)$ and $\theta_{X}$ be as in (4.1). Then for all $j\geq 0$, we have the following identity:
@@
Let $f_{t}\in Aut(M)$ be the flow of ${\rm Re\,}X=(X+\overline{X})/2$. Let $\omega_{t}:=f_{t}^{*}\omega$. Then $\omega_{t}$ is given by $\omega_{t}=\omega+i\partial\bar{\partial}\varphi_{t}$, with $\dot{\varphi_{t}}=if^{*}_{t}\theta_{X}$; see [Wei02, p8].
@@
which completes the proof. ∎
@@
Recall that the Mabuchi functional integrates the Futaki invariant, which does not depend on Kähler metrics. Thus, (4.8) suggests the following proposition. Although this is a special case of [Fut04, Theorem 1.1] (by (4.7)), we include the proof for the reader’s convenience; see also [Tia00b, Remark 3.4].
@@ Appendix A Explicit computations on the generalized Liouville action
Recall the $1$-forms $\gamma^{j}$ defined on $\mathcal{K}_{\omega}$ by (3.39). In this appendix, we show that $\gamma^{2}$ is closed and obtain the second variation formula for the generalized Liouville action $S_{2}$. Note that here we do not assume the polarization of $[\omega]$. We start with some standard identities in Kähler geometry. A good reference for Kähler geometry is [Szé14]. Let $\alpha,\beta$ be $(1,1)$-forms given by $\alpha=i\alpha_{j\bar{k}}dz^{j}\wedge d\bar{z}^{k},\beta=i\beta_{j\bar{k}}dz^{j}\wedge d\bar{z}^{k}$ such that $\alpha_{j\bar{k}},\beta_{j\bar{k}}$ are Hermitian matrices. Then we have
@@
Define $1$-form $\gamma^{2}$ on $\mathcal{K}_{\omega}$ by
$\gamma^{2}_{\varphi}(\psi):=\int_{M}\psi\left(\frac{1}{6}\Delta_{\varphi}S_{\varphi}-\frac{1}{24}\left(|R_{\varphi}|^{2}-4|\operatorname{Ric}_{\varphi}|^{2}+3S_{\varphi}^{2}\right)\right)\frac{\omega_{\varphi}^{n}}{n!}$ (A.5)
for $\psi\in T_{\varphi}\mathcal{K}_{\omega}=C^{\infty}(M,\mathbb{R})$. Then $\gamma^{2}$ is closed.
@@
In this first proof, we prove $\gamma^{2}$ is exact. That is, let $\widetilde{S}:\mathcal{K}_{\omega}\rightarrow\mathbb{R}$ by
@@
Let $\psi\in T_{\varphi}\mathcal{K}_{\omega}$. We will show that $d\widetilde{S}_{\varphi}(\psi)=\left.\frac{d}{dt}\right|_{t=0}\widetilde{S}(\varphi+t\psi)=\gamma^{2}_{\varphi}(\psi)$. It is clear that by virtue of cocyle identity (see Remark 2), we can assume $\varphi=0$ without loss of generality. First note that
@@
In this second proof, we directly prove $\gamma^{2}$ is closed in the spirit of [Mab86]. Let $\psi_{1},\psi_{2}\in T_{\varphi}\mathcal{K}_{\omega}=C^{\infty}(M,\mathbb{R})$. We will show that $d\gamma^{2}_{\varphi}(\psi_{1},\psi_{2})=0$. Note that since
$d\gamma^{2}_{\varphi}(\psi_{1},\psi_{2})=\psi_{1}\cdot\gamma^{2}(\psi_{2})-\psi_{2}\cdot\gamma^{2}(\psi_{1}),$ (A.18)
we only need to show $\psi_{1}\cdot\gamma^{2}(\psi_{2})=\left.\frac{d}{dt}\right|_{t=0}\gamma^{2}_{\varphi+t\psi_{1}}(\psi_{2})$ is symmetric in $\psi_{1}$ and $\psi_{2}$. Assume $\varphi=0$ without loss of generality. First we divide $\gamma^{2}$ into three parts :
$24n!\gamma_{0}(\psi)$ $=\int_{M}4\psi\Delta S\omega^{n}+\int_{M}3\psi\left(|\operatorname{Ric}|^{2}-S^{2}\right)\omega^{n}+\int_{M}\psi\left(|\operatorname{Ric}|^{2}-|R|^{2}\right)\omega^{n}$ (A.19)
@@
$\psi_{1}\cdot\gamma^{2}(\psi_{2})=\left(\text{terms symmetric in $\psi_{1}$ and $\psi_{2}$}\right)$

### v3

@@ 1.2. Relation to physics literature
In the physics literature, the determinantal point process on Riemann surfaces corresponds to the plasma analogy in the quantum Hall effect (QHE). For a physics background on QHE, see [Ton16]. The form of asymptotic expansion of the corresponding partition function was conjectured in [ZW06] and later proved in [Kle14, KMMW17, SY25] for a pure bulk situation. For mathematical results in the presence of an edge, see [Ser24, BKS23, BKSY25] and references therein.
@@ 1.3. Statement of the main results
The following Theorem generalizes formula [Lu04, (4.4)] for all $j\geq 0$. Let $X$ be a holomorphic vector field on $M$ and $\theta_{X}$ be a purely imaginary function satisfying $\iota_{X}\omega=-\bar{\partial}\theta_{X}$.
@@
Let $(M,\omega)$ be a polarized Kähler manifold. Let $X$ and $\theta_{X}$ be given as above. Then for all $j\geq 0$, the following identity holds :
@@ 3. Asymptotic expansion of the DPP partition function
$\gamma^{(j)}_{\varphi}(\psi):=\int_{M}\psi(\Delta_{\varphi}a_{j-1}(\omega_{\varphi})-a_{j}(\omega_{\varphi}))\frac{\omega_{\varphi}^{n}}{n!},$ (3.39)
@@ 4. Holomorphic invariants from Bergman kernel asymptotics and critical points
Let $(M,\omega)$ be a polarized Kähler manifold. Let $X\in\mathfrak{h}(M)$ and $\theta_{X}$ be as in (4.1) and suppose it is purely imaginary. Then for all $j\geq 0$, we have the following identity:
@@
For non-purely imaginary $\theta_{X}$, (4.7) holds with $\theta_{X}$ replaced by $i\,{\rm Im\,}\theta_{X}$ and $\nabla X$ replaced by ${\rm Re\,}\nabla X$.
@@
Assume $\theta_{X}$ to be purely imaginary. Let $f_{t}\in Aut(M)$ be the flow of ${\rm Re\,}X=(X+\overline{X})/2$. Let $\omega_{t}:=f_{t}^{*}\omega$. Then $\omega_{t}$ is given by $\omega_{t}=\omega+i\partial\bar{\partial}\varphi_{t}$, with $\dot{\varphi_{t}}=if^{*}_{t}\theta_{X}$; see [Wei02, p8].
@@
which completes the proof. In the case of non-purely imaginary $\theta_{X}$, the proof is nearly identical to the one given above. ∎
@@
For simplicity, we assume $\theta_{X}$ to be purely imaginary from now on. Recall that the Mabuchi functional integrates the Futaki invariant, which does not depend on Kähler metrics. Thus, (4.8) suggests the following proposition. Although this is a special case of [Fut04, Theorem 1.1] (by (4.7)), we include the proof for the reader’s convenience; see also [Tia00b, Remark 3.4].
@@ Appendix A Explicit computations on the generalized Liouville action
Recall the $1$-forms $\gamma^{(j)}$ defined on $\mathcal{K}_{\omega}$ by (3.39). In this appendix, we show that $\gamma^{(2)}$ is closed and obtain the second variation formula for the generalized Liouville action $S_{2}$. Note that here we do not assume the polarization of $[\omega]$. We start with some standard identities in Kähler geometry. A good reference for Kähler geometry is [Szé14]. Let $\alpha,\beta$ be $(1,1)$-forms given by $\alpha=i\alpha_{j\bar{k}}dz^{j}\wedge d\bar{z}^{k},\beta=i\beta_{j\bar{k}}dz^{j}\wedge d\bar{z}^{k}$ such that $\alpha_{j\bar{k}},\beta_{j\bar{k}}$ are Hermitian matrices. Then we have
@@
Define $1$-form $\gamma^{(2)}$ on $\mathcal{K}_{\omega}$ by
$\gamma^{(2)}_{\varphi}(\psi):=\int_{M}\psi\left(\frac{1}{6}\Delta_{\varphi}S_{\varphi}-\frac{1}{24}\left(|R_{\varphi}|^{2}-4|\operatorname{Ric}_{\varphi}|^{2}+3S_{\varphi}^{2}\right)\right)\frac{\omega_{\varphi}^{n}}{n!}$ (A.5)
for $\psi\in T_{\varphi}\mathcal{K}_{\omega}=C^{\infty}(M,\mathbb{R})$. Then $\gamma^{(2)}$ is closed.
@@
In this first proof, we prove $\gamma^{(2)}$ is exact. That is, let $\widetilde{S}:\mathcal{K}_{\omega}\rightarrow\mathbb{R}$ by
@@
Let $\psi\in T_{\varphi}\mathcal{K}_{\omega}$. We will show that $d\widetilde{S}_{\varphi}(\psi)=\left.\frac{d}{dt}\right|_{t=0}\widetilde{S}(\varphi+t\psi)=\gamma^{(2)}_{\varphi}(\psi)$. It is clear that by virtue of cocyle identity (see Remark 2), we can assume $\varphi=0$ without loss of generality. First note that
@@
In this second proof, we directly prove $\gamma^{(2)}$ is closed in the spirit of [Mab86]. Let $\psi_{1},\psi_{2}\in T_{\varphi}\mathcal{K}_{\omega}=C^{\infty}(M,\mathbb{R})$. We will show that $d\gamma^{(2)}_{\varphi}(\psi_{1},\psi_{2})=0$. Note that since
$d\gamma^{(2)}_{\varphi}(\psi_{1},\psi_{2})=\psi_{1}\cdot\gamma^{(2)}(\psi_{2})-\psi_{2}\cdot\gamma^{(2)}(\psi_{1}),$ (A.18)
we only need to show $\psi_{1}\cdot\gamma^{(2)}(\psi_{2})=\left.\frac{d}{dt}\right|_{t=0}\gamma^{(2)}_{\varphi+t\psi_{1}}(\psi_{2})$ is symmetric in $\psi_{1}$ and $\psi_{2}$. Assume $\varphi=0$ without loss of generality. First we divide $\gamma^{(2)}$ into three parts :
$24n!\gamma^{(2)}_{0}(\psi)$ $=\int_{M}4\psi\Delta S\omega^{n}+\int_{M}3\psi\left(|\operatorname{Ric}|^{2}-S^{2}\right)\omega^{n}+\int_{M}\psi\left(|\operatorname{Ric}|^{2}-|R|^{2}\right)\omega^{n}$ (A.19)
@@
$\psi_{1}\cdot\gamma^{(2)}(\psi_{2})=\left(\text{terms symmetric in $\psi_{1}$ and $\psi_{2}$}\right)$